# Harnessing Earnings Reports for Stock Predictions: A QLoRA-Enhanced LLM Approach

### v1

@@ I Introduction
The rapid growth and development of Artificial Intelligence have created an ideal environment for the emergence of Large Language Models (LLMs), which are transforming the field of natural language processing in many areas[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] and have sparked a great deal of interest in the diverse applications of these models. There are already established works applying advanced modeling techniques, including deep learning [12, 13, 14, 15, 16, 17, 18] and LLMs, in finance. For instance, Zhang et al. [19] fine-tuned an LLM using Retrieval Augmented Generation (RAG) for sentiment analysis, showcasing the use of advanced AI to enhance financial data interpretation. Similarly, Koa et al. [20] introduced the Summarize-Explain-Predict (SEP) framework, which employs self-reflective agents and Proximal Policy Optimization (PPO) to fine-tune LLMs for explainable stock predictions. Additionally, the FinBERT-LSTM model, which combines the financial-domain-focused capabilities of FinBERT with the sequential data processing strength of LSTM, demonstrates superior predictive accuracy in forecasting stock prices based on financial news sentiment analysis over traditional LSTM and DNN models [21]. Complementing this, Chang’s work [22] has shown that integrating advanced Transformer models with dimensionality reduction techniques results in highly efficient classifiers, which are particularly effective in domains such as stock prediction and fraud detection. These advancements highlight the significant role of sophisticated AI in enhancing the accuracy and efficiency of financial market analysis by effectively capturing complex patterns and dependencies within the data.
@@ III-B Instruction Fine Tuning
Instruction fine-tuning is a technique used to enhance the performance of large language models by training them to comprehend and execute specific, structured instructions, especially useful in domains where precise and actionable outputs are needed from complex input data [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. The process involves three key steps: First, creating an instruction-following dataset consisting of paired instructions and their corresponding expected responses, which act as labels for investor actions such as ‘Long’ or ‘Short’. This dataset includes relevant financial data points tailored to reflect real-world scenarios the model will encounter. Second, the model is fine-tuned using this dataset, where it learns to generate the expected responses based on the provided instructions, adapting its pre-existing capabilities to the specific task of making investment predictions [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Finally, the generated outputs are mapped back to the predefined labels to ensure the model’s responses are accurate and applicable to practical financial decision-making settings.
@@ IV-A Model Training
Optimization utilized the adamw_8bit”, an $8$-bit variant of the AdamW optimizer, striking a balance between performance and memory efficiency. Weight decay was set at $0.01$ to regularize the training process and prevent overfitting [56, 57, 58, 59, 60, 61, 62, 63, 64]. The learning rate scheduler was configured to a linear” setting to ensure a steady decrease in the learning rate as training advanced [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83]. Models were capable of processing inputs up to a maximum sequence length of $25,000$, allowing extensive financial texts to be analyzed comprehensively. This structured training approach was designed to enhance models’ predictive accuracy and operational efficiency, capitalizing on their unique architectural strengths.
@@ IV-C1 Evaluation Metrics
To assess the performance of the fine-tuned models, three key metrics were employed: accuracy, weighted F1 score, and the Matthews Correlation Coefficient (MCC). Accuracy measures the overall correctness of predictions, providing a straightforward metric of performance across all classes. The weighted F1 score, which considers both precision and recall, is particularly useful in situations where class imbalance might affect the model’s evaluation. Lastly, the Matthews Correlation Coefficient (MCC) offers a comprehensive measure of the model’s quality, capturing true and false positives and negatives in a single coefficient [84]. This combination of metrics ensures a robust analysis of the models’ predictive capabilities.
@@ V Conclusion and future work
This research demonstrated the efficacy of employing advanced machine learning techniques, specifically instruction-based fine-tuning coupled with QLoRA compression, to enhance the predictive accuracy of stock market forecasts post-earnings announcements. By systematically integrating both ‘base factors’, such as financial metric growth and earnings transcripts, with ‘external factors’, including past week market indices and stock performance alongside analyst grades, this study crafted a comprehensive dataset that enabled the fine-tuned models to capture the multifaceted nature of financial data effectively. The instruction-based fine-tuning process, tailored to the nuanced demands of financial forecasting, significantly augmented the capabilities of our models. This was evidenced by the substantial improvements in performance metrics, particularly when evaluated on the ‘Full’ dataset, which included a broader range of contextual information compared to the ‘Base’ dataset. The LLM, particularly the llama-3-8b-Instruct-4bit model, displayed remarkable performance enhancements, outstripping traditional approaches such as GPT-4 in terms of accuracy, weighted F1, and MCC. Additionally, the effectiveness of fine-tuning transformer-based large language models suggests a promising foundation for future research in various domains. This underscores their potential not only in enhancing financial market predictions but also in extending these techniques to other complex tasks such as credit fraud detection, where similar fine-tuning approaches can develop powerful classification models [22]. Furthermore, the application of QLoRA provided an efficient means of deploying these advanced models on resource-constrained devices, maintaining high levels of predictive accuracy while minimizing operational demands. This aspect of model compression is crucial for real-time financial applications, where timely and effective decision-making is essential.

### v2

@@ I Introduction
The rapid growth and development of Artificial Intelligence have created an ideal environment for the emergence of Large Language Models (LLMs), which are transforming the field of natural language processing in many areas[1, 2, 3] and have sparked a great deal of interest in the diverse applications of these models. There are already established works applying advanced modeling techniques, including deep learning and LLMs, in finance. For instance, Zhang et al. [4] fine-tuned an LLM using Retrieval Augmented Generation (RAG) for sentiment analysis, showcasing the use of advanced AI to enhance financial data interpretation. Similarly, Koa et al. [5] introduced the Summarize-Explain-Predict (SEP) framework, which employs self-reflective agents and Proximal Policy Optimization (PPO) to fine-tune LLMs for explainable stock predictions. Additionally, the FinBERT-LSTM model, which combines the financial-domain-focused capabilities of FinBERT with the sequential data processing strength of LSTM, demonstrates superior predictive accuracy in forecasting stock prices based on financial news sentiment analysis over traditional LSTM and DNN models [6]. These advancements highlight the significant role of sophisticated AI in enhancing the accuracy and efficiency of financial market analysis by effectively capturing complex patterns and dependencies within the data.
@@ III-B Instruction Fine Tuning
Instruction fine-tuning is a technique used to enhance the performance of large language models by training them to comprehend and execute specific, structured instructions, especially useful in domains where precise and actionable outputs are needed from complex input data [7, 8]. The process involves three key steps: First, creating an instruction-following dataset consisting of paired instructions and their corresponding expected responses, which act as labels for investor actions such as ‘Long’ or ‘Short’. This dataset includes relevant financial data points tailored to reflect real-world scenarios the model will encounter. Second, the model is fine-tuned using this dataset, where it learns to generate the expected responses based on the provided instructions, adapting its pre-existing capabilities to the specific task of making investment predictions [9, 10]. Finally, the generated outputs are mapped back to the predefined labels to ensure the model’s responses are accurate and applicable to practical financial decision-making settings.
@@ IV-A Model Training
Optimization utilized the adamw_8bit”, an $8$-bit variant of the AdamW optimizer, striking a balance between performance and memory efficiency. Weight decay was set at $0.01$ to regularize the training process and prevent overfitting [13]. The learning rate scheduler was configured to a linear” setting to ensure a steady decrease in the learning rate as training advanced. Models were capable of processing inputs up to a maximum sequence length of $25,000$, allowing extensive financial texts to be analyzed comprehensively. This structured training approach was designed to enhance models’ predictive accuracy and operational efficiency, capitalizing on their unique architectural strengths.
@@ IV-C1 Evaluation Metrics
To assess the performance of the fine-tuned models, three key metrics were employed: accuracy, weighted F1 score, and the Matthews Correlation Coefficient (MCC). Accuracy measures the overall correctness of predictions, providing a straightforward metric of performance across all classes. The weighted F1 score, which considers both precision and recall, is particularly useful in situations where class imbalance might affect the model’s evaluation. Lastly, the Matthews Correlation Coefficient (MCC) offers a comprehensive measure of the model’s quality, capturing true and false positives and negatives in a single coefficient. This combination of metrics ensures a robust analysis of the models’ predictive capabilities.
@@ V Conclusion and future work
This research demonstrated the efficacy of employing advanced machine learning techniques, specifically instruction-based fine-tuning coupled with QLoRA compression, to enhance the predictive accuracy of stock market forecasts post-earnings announcements. By systematically integrating both ‘base factors’, such as financial metric growth and earnings transcripts, with ‘external factors’, including past week market indices and stock performance alongside analyst grades, this study crafted a comprehensive dataset that enabled the fine-tuned models to capture the multifaceted nature of financial data effectively. The instruction-based fine-tuning process, tailored to the nuanced demands of financial forecasting, significantly augmented the capabilities of our models. This was evidenced by the substantial improvements in performance metrics, particularly when evaluated on the ‘Full’ dataset, which included a broader range of contextual information compared to the ‘Base’ dataset. The LLM, particularly the llama-3-8b-Instruct-4bit model, displayed remarkable performance enhancements, outstripping traditional approaches such as GPT-4 in terms of accuracy, weighted F1, and MCC. Additionally, the effectiveness of fine-tuning transformer-based large language models suggests a promising foundation for future research in various domains. Furthermore, the application of QLoRA provided an efficient means of deploying these advanced models on resource-constrained devices, maintaining high levels of predictive accuracy while minimizing operational demands. This aspect of model compression is crucial for real-time financial applications, where timely and effective decision-making is essential.